# Sequential Conformal Risk Control UQ Uncertainty Quantification

### v1

@@ III-B Sequential Conformal Risk Control with 1+2 parameters
  • Then, two independent steps are carried out in parallel, those of localization and classification: they depend on the first one and are therefore denoted by $L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$, where $\bullet\in\{\mathrm{loc},\mathrm{cls}\}$.
Though not written explicitly in this section, the random loss functions $L^{\mathrm{cnf}}_{i}:\Lambda^{\mathrm{cnf}}\rightarrow\mathbb{R}$ and $L^{\bullet}_{i}:\Lambda^{\mathrm{cnf}}\times\Lambda^{\bullet}\rightarrow\mathbb{R}$ are obtained through task-specific losses and predictions sets defined in Section IV, roughly of the form:888 Though not written explicitly below, the loss $L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$ also depends on a matching between ground-truth and predicted boxes. See Section IV-A for details.
@@
The random loss functions $L^{\mathrm{cnf}}_{i}:\Lambda^{\mathrm{cnf}}\rightarrow\mathbb{R}$ and $L^{\bullet}_{i}:\Lambda^{\mathrm{cnf}}\times\Lambda^{\bullet}\rightarrow\mathbb{R}$ are non-increasing and right-continuous in all of their parameters,999i.e., non-increasing and right-continuous in each of the parameters, all the other parameters being fixed. (This should hold true for every sample path.) and take their values in $\left[0,B^{\mathrm{cnf}}\right]$ and $\left[0,B^{\bullet}\right]$ respectively. Also assume that for any $i\in\{1,\cdots,n\}$, we have $L^{\bullet}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=0$ almost surely.
@@ III-B1 The SeqCRC method
III-B2 Theoretical analysis of SeqCRC
We now state our main theoretical result. It implies a probabilistic guarantee for the localization and classification tasks, and an additional guarantee for the confidence task under an additional assumption on the confidence loss function.
@@ III-B2 Theoretical analysis of SeqCRC
Suppose that Assumptions 1 and 3 hold true. [^10] To be more rigorous, in addition to Assumption 1, we should assume that $(L^{\text{cnf}}_{i},L^{\text{loc}}_{i},L^{\text{cls}}_{i})$ is fully determined (in a measurable way) by $(X_{i},Y_{i})$, as in the examples of Section IV. This implies the following useful property: that the random $3$-tuples $(L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})$ are independent and identically distributed. Let $\alpha^{\mathrm{cnf}}\geq 0$ and $\alpha^{\bullet}\geq\alpha^{\mathrm{cnf}}+\frac{B^{\bullet}}{n+1}$. Then, $\lambda^{\mathrm{cnf}}_{\pm}$ and $\lambda^{\bullet}_{+}$ introduced in (4), (5) and (6) are well defined, and
@@
Under the additional assumption that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i\in\{1,\ldots,n+1\}$,111111 Recall that $\bar{\lambda}^{\mathrm{cnf}}=\max\Lambda^{\mathrm{cnf}}$. we also have
@@
On the monotone loss assumption. Though intuitive, the assumption that the loss functions are monotone can fail in practice for $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$. While it may not appear obvious from the loss definitions in the next section, the matching required for their computation is the root of this issue (see Remarks 1 and 2 for details). This implies that the localization and classification loss functions are typically non-monotone in the $\lambda^{\mathrm{cnf}}$ parameter, even if they are close to monotone in most of our experiments.
To avoid this (small) gap between theory and practice, in Section IV-E, we propose a simple monotonization trick: within (4), (5), and (6) we replace $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$ with tight upper bounds that are provably monotone, as done similarly in [53]. These modified loss functions can be computed on the fly, and preserve the guarantees (7) and (8). Nevertheless, this comes at a cost, as these conservative losses lead to larger parameters $\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ and therefore larger prediction sets, as the gap to monotonicity of the loss increases. This gap should therefore be as small and localized as possible, especially for values below $\alpha$.
@@ IV Application-tailored Prediction Sets & Losses
The general method presented in the previous section can be instantiated with various losses and prediction sets to meet specific practical requirements. This section presents a selection of such sets and losses and compares their features. They are introduced for confidence, localization and classification tasks respectively in Sections IV-B, IV-C and IV-D. Additional losses and prediction sets are discussed in Appendix E. Finally, in Section IV-E we present algorithms with their high-level pseudo-codes both for calibration (where we monotonize irregular losses on the fly, in order to meet the assumptions behind SeqCRC) and for inference.
@@ IV-A Level of Guarantee & Matching
The two settings are strongly connected, and while the former is closer to the usual approach in OD, we focus on the more general image-level one in the core of our article. We argue that the image-level setting allows for more flexibility in the design of losses, and in particular to account for interactions between objects on an image (more details in Appendix A).
@@
Hereon, we assume that every image $x$ produces a matching $\pi_{x}$ that associates each ground-truth $(b_{j},c_{j})$ box/class pair with its nearest prediction $(b_{\pi_{x}(j)},c_{\pi_{x}(j)})$ with respect to a “distance” function $d$ [^12] We use the term “distance” even if $d$ is not necessarily a metric from a mathematical viewpoint.
@@
$d_{\mathrm{mix}}\left((b,c),(\hat{b},\hat{c})\right)=\tau d_{\mathrm{LAC}}(c,\hat{c})+(1-\tau)d_{\mathrm{haus}}(b,\hat{b}).$ (10)
@@
We now address the confidence, localization and classification tasks. Recall that throughout this section we work with a generic example $(x,y)$, which can belong to calibration or test data. The notation $x$ stands for the input image, while $y$ denotes the list (or sequence) of all annotated ground-truth objects on $x$, which are elements of ×$\mathcal{B}\times\mathcal{C}$ (bounding box and class, see Section II-A).
@@ IV-B First Step: Confidence
For the first step introduced in Sec. III-B1, and as (2) in Fig. 1 we define our prediction set and loss as follows.
@@
=$\mathcal{I}=\bigl{\{}k\in\{1,\cdots,N^{\mathrm{nms}}\}:o(x)_{[k]}\geq 1-\lambda^{\mathrm{cnf}}\bigr{\}}\;,$
@@
This loss imposes a strong penalty because having only one less prediction than the number of ground truths yields a loss of $1$. The “box-count-recall” loss is a relaxed version:
@@
These losses do not account for the locations of the predictions and ground truths. We introduce and discuss additional examples of losses in Appendix E.
@@ IV-C Second Step: Localization
For the second step introduced in Sec. III-B1 and illustrated in Fig. 3, recall that by “localization” we refer to how well a bounding box contains the ground-truth bounding box of an object, regardless of its class. Similarly to the previous confidence step, we work with sequences of bounding boxes (multiple objects per image).
@@
Note that, as already done for regression tasks (in, e.g., [50, 63, 54]), we could replace these width and height factors by any notion of local uncertainty to improve the adaptivity of our conformal correction (see Appendix E). It is also possible to obtain corrections that depend on the class associated with the box [23].
@@
Figure 4: Monotonization of a Localization Loss. The plot represents the box-wise recall loss (Eq. (14)) as a function of the confidence parameter $\lambda^{\mathrm{cnf}}$. The various curves (colors) correspond to different fixed values of $\lambda^{\mathrm{loc}}$. The solid lines are the raw losses, and the dotted ones are their monotonized variants. While we could expect raw losses to be non-increasing, they are usually not, due to (partially) diverging goals between the localization loss and the distance $d$ used in the matching algorithm.
See Fig. 4 for an illustration. The reason is that, as $\lambda^{\mathrm{cnf}}$ increases, more and more boxes are added in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$, so that the matching $\pi_{x}$ between true boxes $b_{j}$ and boxes $\hat{b}_{k}$ in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$ changes. This change is guided by the “distance” function $d$ used to compute $\pi_{x}$, which is only vaguely related to the three loss examples above. See Appendix E-D for additional details.
In Section IV-E we will introduce a monotonization trick to fix this issue.
@@ IV-D Classification
The APS variant by [64] consists in retaining the most likely classes too, but in a way that the cumulative softmax probability is above a given threshold. More precisely, one first ranks the softmax scores in decreasing order: let $\kappa_{[1]},\cdots,\kappa_{[K]}\in\mathcal{C}$ be such that $\hat{c}_{k}(\kappa_{[1]})\geq\cdots\geq\hat{c}_{k}(\kappa_{[K]})$. (The order can be different for different values of $k$.) The APS prediction set is then defined by [^13] When $\lambda^{\mathrm{cls}}=1$, we set $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{k}=\mathcal{C}$ by convention.
@@
Let $c_{j}\in\mathcal{C}$ denote the class label of the $j$-th object in the ground truth $y$. We evaluate a binary loss for each true object, by checking whether its class $c_{j}$ belongs to the associated prediction set component $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{\pi_{x}(j)}$. We then aggregate these loss values. In our experiments we compute an average over all true objects, but some alternatives are discussed in Appendix A. The average version reads:
@@ IV-E Algorithm
11:$\lambda^{\mathrm{cnf}}_{+}\leftarrow$ Solve Eq. (4) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, the sets $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}$, $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}$ and $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ , and the losses $L_{i}^{\mathrm{cnf}}$, $L_{i}^{\mathrm{loc}}$, $L_{i}^{\mathrm{cls}}$ using Algorithm 3 (with the monotonization trick)
12:$\lambda^{\mathrm{cnf}}_{-}\leftarrow$ Solve Eq. (5) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, the sets $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}$, $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}$ and $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ , and the losses $L_{i}^{\mathrm{cnf}}$, $L_{i}^{\mathrm{loc}}$, $L_{i}^{\mathrm{cls}}$ using Algorithm 3 (with the monotonization trick)
@@
19:$\lambda^{\mathrm{loc}}_{+}\leftarrow$ Solve Eq. (6) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, $L_{i}^{\mathrm{loc}}$ and $\Gamma^{\mathrm{loc}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}}}$ using Algorithm 4 (with the monotonization trick)
@@
21:$\lambda^{\mathrm{cls}}_{+}\leftarrow$ Solve Eq. (6) using $(X_{i},Y_{i},\hat{Y}_{i})_{i=1}^{n}$, $L_{i}^{\mathrm{cls}}$ and $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}$ using Algorithm 4 (with the monotonization trick)
In Algorithm 1 we present a high-level pseudo-code to compute the parameters $\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cnf}}_{-},\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ defined in (4), (5) and (6), while meeting the monotonicity requirements on the loss functions $L^{\mathrm{loc}}$ and $L^{\mathrm{cls}}$. (Recall from Section IV, Remarks 1 and 2, that these losses are not necessarily monotone in $\lambda^{\mathrm{cnf}}$.) As mentioned in Section III-B, this is achieved via a monotonization trick (e.g., [53]), i.e., by replacing $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$ with tight upper bounds that are provably monotone. More specifically, Algorithm 1 uses two monotonization subroutines: Algorithm 3 for the confidence step, and Algorithm 4 for the localization/classification step, presented in Appendix F.
These two algorithms “monotonize” the loss on the fly, to improve computational and memory efficiency. Due to their complexity and their implementation not being straightforward, more details are given in Appendix F.

### v2

@@ III-B1 Notation and assumptions
  • Then, two independent steps are carried out in parallel: localization and classification. They depend on the first one, and their associated losses are thus denoted by $L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$, where $\bullet\in\{\mathrm{loc},\mathrm{cls}\}$.
Though not written explicitly in this section, the random loss functions $L^{\mathrm{cnf}}_{i}:\Lambda^{\mathrm{cnf}}\rightarrow\mathbb{R}$ and $L^{\bullet}_{i}:\Lambda^{\mathrm{cnf}}\times\Lambda^{\bullet}\rightarrow\mathbb{R}$ are obtained through task-specific losses and predictions sets defined in Section IV, roughly of the form:Though not written explicitly below, the loss $L^{\bullet}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\bullet})$ also depends on a matching between ground-truth and predicted boxes. See Section IV-A for details.
@@
The random loss functions $L^{\mathrm{cnf}}_{i}:\Lambda^{\mathrm{cnf}}\rightarrow\mathbb{R}$ and $L^{\bullet}_{i}:\Lambda^{\mathrm{cnf}}\times\Lambda^{\bullet}\rightarrow\mathbb{R}$ are non-increasing and right-continuous in all of their parameters,i.e., non-increasing and right-continuous in each of the parameters, all the other parameters being fixed. (This should hold true for every sample path.) and take their values in $\left[0,B^{\mathrm{cnf}}\right]$ and $\left[0,B^{\bullet}\right]$ respectively. Also assume that for any $i\in\{1,\cdots,n\}$, we have $L^{\bullet}_{i}(\bar{\lambda}^{\mathrm{cnf}},\bar{\lambda}^{\bullet})=0$ almost surely.
@@ III-B2 The SeqCRC method
III-B3 Theoretical analysis of SeqCRC
We now state our main theoretical result. It implies a probabilistic guarantee for the localization and classification tasks, and an additional guarantee for the confidence task under an additional assumption on the confidence loss function. Recall the placeholder $\bullet\in\{\mathrm{loc},\mathrm{cls}\}$ and that $\bar{\lambda}^{\mathrm{cnf}}=\max\Lambda^{\mathrm{cnf}}$.
@@ III-B3 Theoretical analysis of SeqCRC
Suppose that Assumptions 1 and 3 hold true.To be more rigorous, in addition to Assumption 1, we should assume that $(L^{\text{cnf}}_{i},L^{\text{loc}}_{i},L^{\text{cls}}_{i})$ is fully determined (in a measurable way) by $(X_{i},Y_{i})$, as in the examples of Section IV. This implies the following key property: that the random $3$-tuples $(L^{\text{cnf}}_{1},L^{\text{loc}}_{1},L^{\text{cls}}_{1}),\ldots,(L^{\text{cnf}}_{n+1},L^{\text{loc}}_{n+1},L^{\text{cls}}_{n+1})$ are independent and identically distributed. Let $\alpha^{\mathrm{cnf}}\geq 0$ and $\alpha^{\bullet}\geq\alpha^{\mathrm{cnf}}+\frac{B^{\bullet}}{n+1}$. Then, $\lambda^{\mathrm{cnf}}_{\pm}$ and $\lambda^{\bullet}_{+}$ introduced in (4), (5) and (6) are well defined, and
@@
Under the additional assumption that $L^{\mathrm{cnf}}_{i}(\bar{\lambda}^{\mathrm{cnf}})\leq\alpha^{\mathrm{cnf}}$ almost surely for all $i\in\{1,\ldots,n\}\cup\{\mathrm{test}\}$, we also have
@@
This guarantee is similar to that in Theorem 3 by [25] (independent work, with another targeted application), but our SeqCRC method achieves a finite-sample guarantee with a single data split. In contrast, their approach relies on standard CRC estimators, thus requiring an additional data split to get a finite-sample guarantee for the second stage. Our method avoids this by introducing a corrected estimator, $\lambda^{\bullet}_{+}$, which is based on the optimistic estimator $\lambda^{\mathrm{cnf}}_{-}$ from the first stage.
@@
On the monotone loss assumption. Though intuitive, the assumption that the loss functions are monotone can fail in practice for $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$. While it may not appear obvious from the loss definitions in the next section, the matching required for their computation is the root of this issue (see Remarks 1 and 2 for details). This implies that the localization and classification loss functions are typically non-monotone in the $\lambda^{\mathrm{cnf}}$ parameter.
To avoid this (small) gap between theory and practice, in Section IV-E, we propose a simple monotonization trick: within (4), (5), and (6) we replace $L^{\mathrm{loc}}_{i}$ and $L^{\mathrm{cls}}_{i}$ with the smallest upper bounds that are provably monotone, as done similarly in [53]. These modified loss functions can be computed on the fly, and preserve the guarantees (7) and (8). Nevertheless, this comes at a cost, as these conservative losses lead to larger parameters $\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ and therefore larger prediction sets, as the gap to monotonicity of the loss increases. This gap should therefore be as small and localized as possible, especially for values below $\alpha$.
@@ IV Application-tailored Prediction Sets & Losses
The general method presented in the previous section can be instantiated with various losses and prediction sets to meet specific practical requirements. This section presents a selection of such sets and losses and compares their features. They are introduced for confidence, localization and classification tasks respectively in Sections IV-B, IV-C and IV-D. Additional losses and prediction sets are discussed in Appendix B. Finally, in Section IV-E we present algorithms with their high-level pseudo-codes both for calibration (where we monotonize irregular losses on the fly, in order to meet the assumptions behind SeqCRC) and for inference.
@@ IV-A Level of Guarantee & Matching
The two settings are strongly connected, and while the former is closer to the usual approach in OD, we focus on the more general image-level one in the core of our article. We argue that the image-level setting allows for more flexibility in the design of losses, and in particular to account for interactions between objects on an image (more details in Appendix C).
@@
Hereon, we assume that every image $x$ produces a matching $\pi_{x}$ that associates each ground-truth $(b_{j},c_{j})$ box/class pair with its nearest prediction $(b_{\pi_{x}(j)},c_{\pi_{x}(j)})$ with respect to a “distance” function $d$.We use the term “distance” even if $d$ is not necessarily a metric from a mathematical viewpoint.
@@
[…]
@@
We now address the confidence, localization and classification tasks. Recall that throughout this section we work with a generic example $(x,y)$, which can belong to calibration or test data. The notation $x$ stands for the input image, while $y$ denotes the list (or sequence) of all annotated ground-truth objects on $x$, which are elements of $\mathcal{B}\times\mathcal{C}$ (bounding box and class, see Section II-A).
@@ IV-B First Step: Confidence
For the first step introduced in Sec. III-B2, and as (2) in Fig. 1 we define our prediction set and loss as follows.
@@
$\mathcal{I}=\bigl\{k\in\{1,\cdots,N^{\mathrm{nms}}\}:o(x)_{[k]}\geq 1-\lambda^{\mathrm{cnf}}\bigr\}\;,$
@@
This loss is strict: if the number of predictions is even slightly smaller than the number of ground-truth objects, the penalty is maximal (=1=1). The “box-count-recall” loss is a relaxed version:
@@
These losses do not account for the locations of the predictions and ground truths. We introduce and discuss additional examples of losses in Appendix B.
@@ IV-C Second Step: Localization
For the second step introduced in Sec. III-B2 and illustrated in Fig. 3, recall that by “localization” we refer to how well a bounding box contains the ground-truth bounding box of an object, regardless of its class. Similarly to the previous confidence step, we work with sequences of bounding boxes (multiple objects per image).
@@
Note that, as already done for regression tasks (in, e.g., [50, 63, 54]), we could replace these width and height factors by any notion of local uncertainty to improve the adaptivity of our conformal correction (see Appendix B). It is also possible to obtain corrections that depend on the class associated with the box [24].
@@
Figure 4: Monotonization of the Localization Loss. The plot represents the box-wise recall risk on calibration data as a function of the confidence parameter $\lambda^{\mathrm{cnf}}$, that is, the box-wise recall loss (Eq. (14)) averaged over the calibration set. The various colors (blue, green, red, cyan, magenta) correspond to different fixed values of $\lambda^{\mathrm{loc}}$ (0, 125, 250, 375, 500). The solid lines are the raw risks, and the dotted ones are the risks after loss monotonization (cf. Section IV-E). While we could expect raw losses (and thus raw risks) to be non-increasing, they are usually not, due to (partially) diverging goals between the localization loss and the distance $d$ used in the matching algorithm.
See Fig. 4 for an illustration. The reason is that, as $\lambda^{\mathrm{cnf}}$ increases, more and more boxes are added in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$, so that the matching $\pi_{x}$ between true boxes $b_{j}$ and boxes $\hat{b}_{k}$ in $\Gamma^{\mathrm{cnf}}_{\lambda^{\mathrm{cnf}}}(x)$ changes. This change is guided by the “distance” function $d$ used to compute $\pi_{x}$, which is only vaguely related to the three loss examples above.
In Section IV-E and Appendix D we will introduce a monotonization trick to fix this issue.
@@ IV-D Classification
The APS variant by [64] consists in retaining the most likely classes too, but in a way that the cumulative softmax probability is above a given threshold. More precisely, one first ranks the softmax scores in decreasing order: let $\kappa_{[1]},\cdots,\kappa_{[K]}\in\mathcal{C}$ be such that $\hat{c}_{k}(\kappa_{[1]})\geq\cdots\geq\hat{c}_{k}(\kappa_{[K]})$. (The order can be different for different values of $k$.) The APS prediction set is then defined byWhen $\lambda^{\mathrm{cls}}=1$, we set $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{k}=\mathcal{C}$ by convention.
@@
Let $c_{j}\in\mathcal{C}$ denote the class label of the $j$-th object in the ground truth $y$. We evaluate a binary loss for each true object, by checking whether its class $c_{j}$ belongs to the associated prediction set component $\Gamma^{\mathrm{cls}}_{\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}}}(x)_{\pi_{x}(j)}$. We then aggregate these loss values. In our experiments we compute an average over all true objects, but some alternatives are discussed in Appendix C. The average version reads:
@@ IV-E Algorithm
[…]
@@
[…]
@@
[…]
In Algorithm 1 we present a high-level pseudo-code to compute the parameters $\lambda^{\mathrm{cnf}}_{+},\lambda^{\mathrm{cnf}}_{-},\lambda^{\mathrm{loc}}_{+},\lambda^{\mathrm{cls}}_{+}$ defined in (4), (5) and (6), while meeting the monotonicity requirements on the loss functions $L^{\mathrm{loc}}$ and $L^{\mathrm{cls}}$. (Recall from Section IV, Remarks 1 and 2, that these losses are not necessarily monotone in $\lambda^{\mathrm{cnf}}$.) As mentioned in Section III-B, this is achieved via a monotonization trick (e.g., [53]), i.e., by replacing $L^{\mathrm{loc}}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{loc}})$ and $L^{\mathrm{cls}}_{i}(\lambda^{\mathrm{cnf}},\lambda^{\mathrm{cls}})$ with the smallest upper bounds that are provably monotone, namely, $\sup_{\lambda^{\prime}\geq\lambda^{\mathrm{cnf}}}L^{\mathrm{loc}}_{i}(\lambda^{\prime},\lambda^{\mathrm{loc}})$ and $\sup_{\lambda^{\prime}\geq\lambda^{\mathrm{cnf}}}L^{\mathrm{cls}}_{i}(\lambda^{\prime},\lambda^{\mathrm{cls}})$ respectively.
More specifically, Algorithm 1 uses two optimization subroutines detailed in Appendix D: Algorithm 3 for the confidence step, and Algorithm 4 for the localization/classification step. These two algorithms “monotonize” the underlying losses on the fly (while optimizing the risk), to improve computational and memory efficiency.